A 0.053 M solution of an acid (HA) has a pH of 4.63. What is the pKa?
Answer #1

bH=4.63 pH = -log (ht] [H+] = lo - H [ht] = 10-4.63 2.358 10 5 M ICE Table concentration (M) I HA (99) H+ (aq) + A-() 0.053 с